40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 29 17:10:09 2021, Anchal, Summary, CDS, c1teststand problems summary c1teststand_issues_summary.pdf
    Reply  Thu Sep 30 14:09:37 2021, Anchal, Summary, CDS, New way to ssh into c1teststand 
    Reply  Mon Oct 4 11:05:44 2021, Anchal, Summary, CDS, c1teststand problems summary 
       Reply  Mon Oct 4 18:00:16 2021, Koji, Summary, CDS, c1teststand problems summary 
          Reply  Tue Oct 5 17:58:52 2021, Anchal, Summary, CDS, c1teststand problems summary 
       Reply  Tue Oct 5 18:00:53 2021, Anchal, Summary, CDS, c1teststand time synchronization working now 
          Reply  Mon Oct 11 17:31:25 2021, Anchal, Summary, CDS, Fixed mounting of mx devices in fb. daqd_dc is running now. 
             Reply  Mon Oct 11 18:29:35 2021, Anchal, Summary, CDS, Moving forward? 
                Reply  Tue Oct 12 17:20:12 2021, Anchal, Summary, CDS, Connected c1sus2 to martian network 
                   Reply  Tue Oct 12 23:42:56 2021, Koji, Summary, CDS, Connected c1sus2 to martian network 
                      Reply  Wed Oct 13 11:25:14 2021, Anchal, Summary, CDS, Ran c1sus2 models in martian CDS. All good! CDS_screens_running.png
                         Reply  Tue Oct 19 18:20:33 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test data3_Plots.pdfdata2_Plots.pdf
                            Reply  Tue Oct 19 23:43:09 2021, Koji, Summary, CDS, c1sus2 DAC to ADC test P_20211019_224433.jpgP_20211019_224122.jpgP_20211019_224400.jpgP_20211019_224411.jpg
                               Reply  Wed Oct 20 11:48:27 2021, Anchal, Summary, CDS, Power supple configured correctly. 
                               Reply  Tue Oct 26 18:24:00 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test data2_Plots.pdfdata3_Plots.pdf
             Reply  Tue Oct 12 17:10:56 2021, Anchal, Summary, CDS, Some more information 
Message ID: 16417     Entry time: Wed Oct 20 11:48:27 2021     In reply to: 16415
Author: Anchal 
Type: Summary 
Category: CDS 
Subject: Power supple configured correctly. 

This was horrible! That's my bad, I should have checked the configuration before assuming that it is right.

I fixed the power supply configuration. Now the strip has two rails of +/- 18V and the GND is referenced to power supply earth GND.

Ian should redo the tests.

ELOG V3.1.3-